OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (anonymous):

A) f(x) = x^4 - 8x^3 - 12x^2 + 400x - 1599 B) f(x) = x^4 - 200x^2 + 800x - 1599 C) f(x) = x^4 - 98x^2 + 800x - 1599 D) f(x) = x^4 - 8x^3 + 12x^2 - 400x + 1599

OpenStudy (phi):

If a polynomial with real coefficients has complex roots, they *always* come in complex conjugate pairs. In other words, if one of the roots is a+ b*i (where i = sqr(-1) ), there is also the root a - b*i in your case, you have a 4th root (the complex conjugate of 5 + 4i ) = 5 - 4i your polynomial in factored form is ( x - (5+4i)) (x-(5-4i)) (x-3) (x - -13) = ( x - (5+4i)) (x-(5-4i)) (x-3) (x +13)

OpenStudy (anonymous):

How did you do that last step @phi ?

OpenStudy (anonymous):

the last step is to find a polynomial with zero of \(5+4i\) which you can do by working backwards \[x=5+4i\\ x-5=4i\\ (x-5)^2=(4i)^2=-16\\ x^2-10x+25=-16\\ x^2-10x+41=0\]

OpenStudy (anonymous):

or you can know that if \(a+bi\) is a zero of a quadratic, the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

Oh, okay.. Sooo.. I'm real slow right now. I don't want the answer, but I still can't quite understand this...

OpenStudy (phi):

Using Sat's answer, you have \[ ( x - (5+4i)) (x-(5-4i)) (x-3) (x +13) \\ = (x^2-10x+41) (x-3) (x +13) \] To finish, see https://www.khanacademy.org/math/algebra/multiplying-factoring-expression/multiplying_polynomials/v/multiplying-polynomials and the next 2 videos. See if you can do this problem.

OpenStudy (anonymous):

This is going to sound real, real dumb.. How do you multiply three's? I forgot the term.. I am super slow today. Example: \[(x^2-10x+41) (x^2+10x-39)\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

it is sort of like multiplying 123 * 456 by hand, if you learned how to do that